OpenStudy (anonymous):

sin A=15/7 and sin B=3/5. what is the value of cos C?

OpenStudy (mertsj):

Not enough information.

OpenStudy (anonymous):

the information just like that. in triangle ABC known sin A=15/7 and sin B=3/5

OpenStudy (mertsj):

Well now, that is a whole different thing. We now know that the angles are in the same triangle.

OpenStudy (mertsj):

And also, something else is wrong. The sin of an angle cannot be greater than 1 so sin A=15/7 is a non starter.

OpenStudy (anonymous):

i know, it using comparation. there are 3 different triangle. thanks for reply :)

OpenStudy (agent0smith):

lol what the hell. First there's one triangle, now there's three? Just post the entire question, word for word, diagrams included if there are any.